Deployment team

→ Europe/Zurich
EVO - GridPP Deployment team meeting

EVO - GridPP Deployment team meeting

Jeremy Coles
Description
- This is the weekly DTEAM meeting - The intention is to run the meeting in EVO: http://evo.caltech.edu/evoGate/. Join the meeting in the GridPP Community area - - The phone bridge number is +41 22 76 71400. Th phone bridge ID is 124366 with code: 4880.
DTeam - 26 Feb 2008 Attendees ========= Derek Ross Frederic Brochu Andrew Elwell Jeremy Coles Duncan Rand Pete Gronbech Experiments problems/issues =========================== LHCB ---- - CMS --- - CMS functioning well ATLAS ----- CCRC 08 - moved from FDR 1 to CCRC08, testing new version of DDM - went fine. Running reprocess at full speed this week. Data being subscribed to T2. Using srm v2 endpoints already set up but more welcome, testing Sheffield and Brunel at the moment. Cleaning up files from Brunel to make more space. 20TB of space filled up at T1, but only a small fraction going to T2s (but >200GB) CCRC 08 data will be deleted to make space for M6 Cosmics in March. Satisfied with UK Cloud behaviour so far in CCRC08. POSIX I/O use at T2? Other ----- - Site Status ----------- Accounting Enforcement IC-HEP - Investigating UCL-Central - Effectively in maintenance, focusing on new cluster UCL-HEP - Investigating last week, moving machine rooms this week LIV-HEP ECDF RALPP - Investigating - possible new CE problems? SAM QMUL UCL-HEP GLasgow - upgrade last week after CE issue on Wednesday, set user maximum jobs to 500 (from 1000) not sure if this was the original cause though. Some information system issues too Cambridge SL ATLAS Brunel - in maintenance QMUL UCL-HEP Manchester Edinburgh ROC Update ========== PPS being reorganised, PPS site requested to complete the service inventory information PPS Update 19 released DECH reported issued with WMS input sandboxes, jobs go into running state but sandbox does not arrive on WN (See GGUS 33136) Pre-release of SAM web services. Nagios plugin uses it. Ticket Status Importer may not be fixed for a while Site Issues =========== sgm/prod accounts Upgrading to SL4 End of support for certain gLite 3.0 services. See e-mail on DTEAM list, Smaller sites find it harder to move, due to less manpower POSIX I/O Access - Greig is dealing with this. Networking Update ================= Lancaster OPN issue CMS issues appear resolved, but need to confirm Anything for future discussions? Actions Review ============== AOB === Altas Jamboree WLCG Collaboration meeting - about 20 requests to go. CCRC F2F 4th march , Derek giving talk about CCRC site issues Glasgow - GSTAT issues
There are minutes attached to this event. Show them.
    • 11:00 → 11:20
      Experiment problems/issues 20m
      Review of weekly issues by experiment/VO - LHCb - CMS - ATLAS - Other SITE status ************* - Accounting: http://www3.egee.cesga.es/acctenfor/nodata.php -- IC-HEP (9); UCL-CENTRAL (SD); UCL-HEP (SD); LIV-HEP (23); ECDF and RALPP (about to fail) - SAM: QMUL; UCL-HEP; Glasgow; Cambridge all have recent problems - SL ATLAS: Brunel (one cluster); QMUL; UCL-HEP; Manchester and Edinburgh appear to have problems recently.
    • 11:20 → 11:35
      ROC update 15m
      ops meeting ************************* - PPS is being reorganised to aid pre-deployment testing and be more suitable for the LHC VOs. PPS sites are requested to review and complete the Service Inventory information at www.cern.ch/pps/index.php?dir=./site/ - Glite 3.1.0 PPS Update 19 was released to production and it is now in pre-deployment testing * WN 3.1 for sl4 64bits * glite-LSF_utils * lcg-vomscerts-4.8.0 adds next cert for biomed + egeode * new version of lcg-ManageVOTag fixing bug #31848 Release notes in: https://twiki.cern.ch/twiki/bin/view/EGEE/PPSReleaseNotes_310_PPS_Update19 - A new update, gLite3.1.0 PPS Update20 is in preparation. This Update will introduce the MONBOX on the 3.1 baseline (for SLC4) - DECH have seen problems related to the size of the WMS input sand boxes. Have we observered any problems? The WMS is configured to accept up to 100MB. Jobs go in running state but the input sand box does not arrive on the WN. GGUS ticket: 33136. - There is a new pre-release available of the SAM web services. See https://twiki.cern.ch/twiki/bin/view/LCG/SamReleaseActivity#SAM_Validation. Ticket status *************** - See more information link
      more information
    • 11:35 → 11:45
      Site issues 10m
      - Configuration of sgm/prod accounts - Glasgow CE load related problem due to gLite-WMS - Upgrading to SL4 WNs (13/22 at SL4) - End of support for certain gLite 3.0 services - posix i/o access (ATLAS need this for default Ganga submissions)
    • 11:45 → 11:55
      Networking update 10m
      - Any news in this areas (such as improvements in data access) - Problems being seen in CCRC - ...
    • 11:55 → 12:05
      Actions review 10m
    • 12:05 → 12:10
      AOB 5m